# Neural Force Field: Learning Generalized Physical Representation from a Few Examples

### v1

@@ Abstract
Physical reasoning is a remarkable human ability that enables rapid learning and generalization from limited experience. Current AI models, despite extensive training, still struggle to achieve similar generalization, especially in Out-of-distribution (OOD) settings. This limitation stems from their inability to abstract core physical principles from observations. A key challenge is developing representations that can efficiently learn and generalize physical dynamics from minimal data. Here we present Neural Force Field (NFF), a modeling framework built on Neural Ordinary Differential Equation (NODE) that learns interpretable force field representations which can be efficiently integrated through an Ordinary Differential Equation (ODE) solver to predict object trajectories. Unlike existing approaches that rely on high-dimensional latent spaces, NFF captures fundamental physical concepts such as gravity, support, and collision in an interpretable manner. Experiments on two challenging physical reasoning tasks demonstrate that NFF, trained with only a few examples, achieves strong generalization to unseen scenarios. This physics-grounded representation enables efficient forward-backward planning and rapid adaptation through interactive refinement. Our work suggests that incorporating physics-inspired representations into learning systems can help bridge the gap between artificial and human physical reasoning capabilities.
Machine Learning, ICML
@@ 1 Introduction
Physical reasoning, the ability to understand and predict how objects interact in the physical world, is fundamental to both human intelligence and artificial systems (Spelke, 2022). This capability underlies crucial applications ranging from robotics to scientific discovery, making it a central challenge in AI research (Lake et al., 2017). One of the remarkable aspects of human cognitive capabilities is the ability to rapidly learn from limited examples (Kim et al., 2020; Jiang et al., 2022; Lake & Baroni, 2023; Zhang et al., 2024), especially evident in intuitive physics (Kubricht et al., 2017; Bear et al., 2021). Humans can quickly abstract core physical principles after observing limited physical phenomena, enabling them to predict complex dynamics and interact with novel environments (Spelke & Kinzler, 2007; Battaglia et al., 2013; Bonawitz et al., 2019; Xu et al., 2021).
In contrast, current AI systems face significant limitations in physical reasoning. Despite being trained on gigantic datasets, these models still struggle to achieve human-level generalization, particularly in Out-of-distribution (OOD) settings (Lake et al., 2017; Zhu et al., 2020). The core issue lies in their tendency to overfit observed trajectories rather than capturing inherent physical principles, severely limiting their ability to compose known knowledge and predict outcomes in novel contexts (Qi et al., 2021; Li et al., 2022; Wu et al., 2023). This stark contrast between human capabilities and current model limitations has motivated the search for new approaches that can learn generalizable physical representations from minimal data.
@@
To address these challenges, we introduce Neural Force Field (NFF), a physical reasoning framework building upon Neural Ordinary Differential Equation (NODE) for efficient interactive learning and reasoning. At its core, NFF employs a neural network to learn dynamic latent force fields from external interventions and object interactions. These predicted forces are then integrated through an ODE solver to compute explicit physical variables such as velocity and displacement, producing interpretable results that align with established physical principles.
Our framework offers three advantages. First, by representing physical interactions in low-dimensional force fields, NFF can rapidly learn fundamental physical concepts from just a few training examples. Second, due to the ODE-grounded dynamic graph, NFF can effectively generalize to OOD scenarios. Third, the integration of forces through an ODE solver enables fine-grained physical interactions, supporting precise modeling of collisions, gravity effects, and real-time planning.
We validate our approach on two challenging physical reasoning benchmarks: I-PHYRE (Li et al., 2023) and N-body problems (Newton, 1833). These tasks feature complex dynamics ranging from short-range forces (collision and sliding) to long-range forces (gravity), providing a comprehensive test of our framework’s capabilities. Our experiments demonstrate that NFF not only learns dynamics efficiently by abstracting physical interactions into force fields but also achieves strong generalization in both within-scenario and cross-scenario settings. Moreover, the framework’s physics-based representation enables effective forward and backward planning in goal-directed tasks, consistently outperforming existing approaches such as Interaction Network (IN) and transformer-based methods.
@@ 2 Related work
Physical reasoning  Research in physical reasoning has progressed along two main trajectories: passive observation and interactive platforms. The passive observation approach, exemplified by the Violation-of-Expectation (VoE) paradigm (Spelke et al., 1992), evaluates physical understanding by measuring agents’ ability to detect violations of intuitive physics principles (ee Baillargeon, 1987; Hespos & Baillargeon, 2001; Dai et al., 2023). While this approach has provided valuable insights into basic physical comprehension, it is limited by its inability to assess active interventions and complex reasoning.
@@
To enable more comprehensive evaluation, interactive platforms such as PHYRE (Bakhtin et al., 2019), the virtual tools game (Allen et al., 2020), and I-PHYRE (Li et al., 2023) have emerged. These environments require agents to actively manipulate objects to achieve specific goals, testing not only prediction capabilities but also planning and reasoning skills. However, existing approaches in these platforms often struggle with generalization across diverse scenarios, particularly in few-shot settings where limited training data is available (Qi et al., 2021; Li et al., 2023).
@@
Current physical reasoning models face two primary challenges: the need for extensive training data and limited cross-scenario transferability. While some methods achieve strong performance within specific scenarios (Allen et al., 2020), they often fail to generalize their understanding to novel situations, falling short of human-level reasoning capabilities (Kang et al., 2024). Our work addresses these limitations by introducing a framework that unifies passive observation and interactive learning through dynamic force fields. The NFF approach enables few-shot learning of physical principles while supporting active reasoning through its interpretable force-based representation, facilitating both accurate prediction and effective intervention planning across diverse physical scenarios.
Figure 1: Comparison between traditional interaction modeling and NFF. (a) Traditional methods encode physical interactions as high-dimensional latent vectors and predict future states through learned transitions between these vectors. This approach requires extensive training data and often leads to overfitting, limiting generalization to novel scenarios. (b) Our NFF framework represents interactions as learnable force fields predicted from the dynamic object graphs, which are integrated through a differentiable ODE solver to predict trajectories. NFF enables few-shot learning of various interaction types from limited interventions while maintaining strong generalization capabilities. The framework consists of four key operators: E (encoder network for processing physical scenes), D (decoder network for state reconstruction), F (force field predictor), and $\int$ (numerical integrator for computing trajectories).
@@
Dynamic prediction  The prediction of physical dynamics, fundamental to intuitive physics, has evolved along two methodological branches: discrete and continuous-time approaches. Discrete methods typically combine recurrent architectures with GNNs (Battaglia et al., 2016; Qi et al., 2021) or transformer modules (Wu et al., 2023) for modeling object interactions, while leveraging convolutional operators for processing pixel-based information (Shi et al., 2015; Wang et al., 2022). Despite their flexibility in handling various interaction types, these methods often struggle with two key limitations: difficulty in extracting robust physical representations and susceptibility to error accumulation over extended time horizons.
@@
Continuous-time methods address these temporal consistency issues by incorporating physical inductive biases via explicit modeling of state derivatives within dynamical systems (Chen et al., 2018; Greydanus et al., 2019; Zhong et al., 2020; Cranmer et al., 2020; Norcliffe et al., 2020). While this approach improves long-term prediction stability, current systems typically assume energy-conservative systems with simplified dynamics. More importantly, these methods face significant challenges in few-shot learning scenarios and struggle with cross-scenario generalization (Chen et al., 2020), often requiring specific physical priors and grammars to achieve meaningful transfer (Xu et al., 2021).
@@
Our work advances continuous-time methods by introducing a more general physical representation framework via learnable force fields. This approach enables robust cross-scenario generalization while handling complex non-conservative energy systems. By combining the stability advantages of continuous-time modeling with flexible force field representations, NFF achieves both accurate long-term predictions and strong generalization capabilities without requiring explicit physical priors.
@@ 3.1 The Neural Force Field (NFF) representation
Traditional approaches to modeling physical interactions typically rely on implicit representations through hidden vectors to describe object interactions. These methods use neural networks to learn state transition functions that map current scene features to future states. However, this purely data-driven approach cannot guarantee physically grounded representations, leading to poor generalization despite intensive training. We introduce NFF, which addresses these limitations by learning physics-grounded, generalizable representations through explicit force field modeling; see Figure 1 for a comparison.
@@
Prediction force field  The core of NFF is built on the physical concept of a force field—a vector field that determines the force experienced by any query object based on its state $\mathbf{z}^{q}$. For a scene with $N$ objects, we model the force field $\mathbf{F}(\cdot)$ at time $t$ using a neural network operating on a relation graph $\mathcal{G}$. The graph contains object nodes $V=\{\mathbf{z}^{0}(t),\mathbf{z}^{1}(t),...,\mathbf{z}^{N-1}(t)\}$. Following neural operator methods (Lu et al., 2021) and graph neural models (Battaglia et al., 2018), we formulate the force field function as:
$\mathbf{F}(\mathbf{z}^{q}(t))=\sum_{i\in\mathcal{G}(q)}\mathbf{W}\left(f_{\theta}(\mathbf{z}^{i}(t))\cdot f_{\phi}(\mathbf{z}^{q}(t))\right)+\mathbf{b},$ (1)
@@
where $\mathcal{G}(q)$ represents the neighbors of query $q$, $f_{\theta}$ and $f_{\phi}$ are neural networks with parameters $\theta$ and $\phi$, and $\mathbf{W}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{force}}},\mathbf{b}\in\mathbb{R}^{d_{\text{force}}}$ map hidden features to low-dimensional forces. The state vector $\mathbf{z}$ incorporates zero-order (position, angle) and first-order (velocity, angular velocity) states. This representation naturally handles various physical interactions including collision, friction, spring forces, and rotation.
@@
Computing trajectory with ODE Rather than relying on neural decoders, NFF employs a second-order ODE integrator to compute object trajectories from the learned force field. This approach ensures physically consistent trajectories governed by fundamental principles. The dynamic change of motion for a query state follow:
$\frac{d^{2}x^{q}(t)}{dt^{2}}=\frac{dv^{q}(t)}{dt}\propto\mathbf{F}(\mathbf{z}^{q}(t)).$ (2)
@@
While object mass theoretically affects acceleration, these mass constants are absorbed into the learned $F(\cdot)$.
@@
Solving ODEs We solve the ODE system using numerical integration methods such as Runge-Kutta:
$\mathbf{z}^{q}(t)=\text{ODESolve}\left(\mathbf{z}^{q}(0),\mathbf{F},0,t\right),$ (3)
[…]
@@
The NFF framework offers three key advantages over existing approaches. First, it enables few-shot learning of force fields from minimal examples—even single trajectories—through its compact, physically-grounded representation $\mathbf{F}(\mathbf{z}^{q}(t))$. Second, by representing interactions as composable force fields, NFF naturally generalizes to novel scenarios with varying object configurations, interaction types, and time horizons. As shown in Equation 1, this compositionality enables transfer of learned force patterns across physical scenarios through simple force summation. Third, the high-precision integration decoder, formulated in Equations 3 and 4, enables both fine-grained modeling of continuous-time interactions and efficient optimization for planning tasks through its invertible information flow. This precision supports various applications, from determining initial conditions for desired outcomes to designing targeted force fields for specific behaviors. The effectiveness of these three components—low-dimensional representation, ODE grounding, and precise integration—is systematically validated through ablation studies in Section 4.5.
@@
Training NFF The framework employs autoregressive prediction, using previously predicted states to generate current states. Network optimization minimizes the Mean Squared Error (MSE) loss between predictions and ground truth. To promote learning of local, generalizable dynamics rather than global patterns, we segment long trajectories into smaller units during training. Our experiments demonstrate that this approach significantly improves convergence.
(a) I-PHYRE sim (Li et al., 2023)
(b) N-body sim (Rein & Liu, 2012)
(c) Falling
(d) Sliding
(e) Collision
(f) Spring
(g) Visualization of learned physical concepts
(h) Gravitational field ground truth
(i) Learned gravitational field
Figure 2: Visualization of learned physical dynamics and force fields. (a) I-PHYRE simulation environment demonstrating complex rigid body interactions. (b) N-body gravitational simulation showing orbital trajectories. (c) Examples of learned physical concepts demonstrated through trajectory predictions: falling under gravity, sliding with friction, collision with momentum transfer, and spring-like elastic interactions. (d-e) Comparison between ground truth and learned gravitational field distributions shows accurate force field reconstruction. Our NFF successfully captures these fundamental physical behaviors from few-shot learning.
(a) Initial configuration
(b) Ground truth trajectory
(c) Our NFF prediction
(d) SlotFormer (Wu et al., 2023)
(e) IN (Battaglia et al., 2016)
Figure 3: Trajectory predictions on unseen scenarios after few-shot learning. We compare model predictions across four physical scenarios: (i) A seesaw mechanism launching a ball over a wall, demonstrating complex contact dynamics and energy transfer, (ii) A path-building puzzle where white blocks can be eliminated during interaction, requiring the ball to roll across gaps, (iii) A dual-comet gravitational slingshot shows orbital deflection around a massive celestial body, and (iv) An eight-body gravitational system emulating solar system exhibits complex orbital dynamics around a central mass (i.e., Sun). Light trails indicate object trajectories over time. Notably, our NFF predictions closely match the ground truth behaviors across these diverse scenarios, from rigid body interactions to gravitational dynamics. Additional visualizations are provided in Appendix F.

### v4

# Neural Force Field: Few-shot Learning of Generalized Physical Reasoning

@@ Abstract
Physical reasoning is a remarkable human ability that enables rapid learning and generalization from limited experience. Current AI models, despite extensive training, still struggle to achieve similar generalization, especially in Out-of-distribution (OOD) settings. This limitation stems from their inability to abstract core physical principles from observations. A key challenge is developing representations that can efficiently learn and generalize physical dynamics from minimal data. Here we present Neural Force Field (NFF), a framework extending Neural Ordinary Differential Equation (NODE) to learn complex object interactions through force field representations, which can be efficiently integrated through an Ordinary Differential Equation (ODE) solver to predict object trajectories. Unlike existing approaches that rely on discrete latent spaces, NFF captures fundamental physical concepts such as gravity, support, and collision in continuous explicit force fields. Experiments on three challenging physical reasoning tasks demonstrate that NFF, trained with only a few examples, achieves strong generalization to unseen scenarios. This physics-grounded representation enables efficient forward-backward planning and rapid adaptation through interactive refinement. Our work suggests that incorporating physics-inspired representations into learning systems can help bridge the gap between artificial and human physical reasoning capabilities.
@@ 1 Introduction
Physical reasoning, the ability to understand and predict how objects interact in the physical world, is fundamental to both human intelligence and artificial systems (Spelke, , 2022). This capability underlies crucial applications ranging from robotics to scientific discovery, making it a central challenge in AI research (Lake et al., , 2017). One of the remarkable aspects of human cognitive capabilities is the ability to rapidly learn from limited examples (Kim et al., , 2020; Jiang et al., , 2022; Lake and Baroni, , 2023; Zhang et al., , 2024), especially evident in intuitive physics (Kubricht et al., , 2017; Bear et al., , 2021). Humans can quickly abstract core physical principles after observing limited physical phenomena, enabling them to predict complex dynamics and interact with novel environments (Spelke and Kinzler, , 2007; Battaglia et al., , 2013; Bonawitz et al., , 2019; Xu et al., , 2021).
In contrast, current AI systems face significant limitations in physical reasoning. Despite being trained on gigantic datasets, these models still struggle to achieve human-level generalization, particularly in Out-of-distribution (OOD) settings (Lake et al., , 2017; Zhu et al., , 2020). The core issue lies in their tendency to overfit observed trajectories rather than capturing inherent physical principles, severely limiting their ability to compose known knowledge and predict outcomes in novel contexts (Qi et al., , 2021; Li et al., , 2022; Wu et al., , 2023). This stark contrast between human capabilities and current model limitations has motivated the search for new approaches that can learn generalizable physical representations from minimal data.
@@
To address these challenges, we introduce Neural Force Field (NFF), a neural function map from objects to forces acting on continuous Euclidean coordinates for efficient interactive learning and reasoning. At its core, NFF employs a neural operator to learn dynamic force fields from external interventions and object interactions. These predicted forces are then integrated through an ODE solver to compute explicit physical variables such as velocity and displacement, producing interpretable results that align with established physical principles. Please refer to Fig.˜1 for details.
Our framework offers three advantages. First, by representing physical interactions in low-dimensional neural operator force fields, NFF can rapidly learn fundamental physical concepts from just a few training examples. Second, due to the ODE-grounded dynamic graph, NFF can effectively generalize to OOD scenarios. Third, the integration of forces through an ODE solver enables fine-grained physical interactions, supporting precise modeling of collisions and gravity effects.
We evaluate our method using two state-based physical reasoning benchmarks—I-PHYRE (Li et al., , 2023) and N-body problems (Newton, , 1833)—as well as one vision-based benchmark, PHYRE (Bakhtin et al., , 2019). These tasks feature complex rigid-body dynamics ranging from short-range forces (collision and sliding) to long-range forces (gravity). Our experiments demonstrate that NFF not only learns dynamics efficiently by abstracting physical interactions into force fields but also achieves strong generalization in both within-scenario and cross-scenario settings. Moreover, the framework’s physics-based representation enables effective forward and backward planning in goal-directed tasks, consistently outperforming existing GNN-based methods such as Interaction Network (IN) (Battaglia et al., , 2016; Qi et al., , 2021) and transformer-based methods (Wu et al., , 2023).
Figure 1: The framework of NFF. (a) NFF models complex physical interactions by constructing dynamic interaction graphs from scenes and performing continuous integration on force fields. (b) The force fields are inferred by a neural operator that takes object interactions as input. (c) After learning from a few examples, NFF can capture various physical concepts represented by force fields and generalize to unseen OOD scenarios.
@@ 2 Related work
Physical reasoning  Research in physical reasoning has progressed along two main trajectories: passive observation and interactive platforms. The passive observation approach, exemplified by the Violation-of-Expectation (VoE) paradigm (Spelke et al., , 1992), evaluates physical understanding by measuring agents’ ability to detect violations of physics principles (Baillargeon, , 1987; Hespos and Baillargeon, , 2001; Dai et al., , 2023). While this approach has provided valuable insights into basic physical comprehension, it is limited in assessing active interventions and complex reasoning.
@@
To enable more comprehensive evaluation, interactive platforms such as PHYRE (Bakhtin et al., , 2019), the virtual tools game (Allen et al., , 2020), and I-PHYRE (Li et al., , 2023) have emerged. These environments require agents to actively manipulate objects to achieve specific goals, testing not only prediction capabilities but also planning and reasoning skills. However, existing approaches in these platforms often struggle with generalization across diverse scenarios, particularly in few-shot settings where limited training data is available (Qi et al., , 2021; Li et al., , 2023).
@@
Current physical reasoning models face two primary challenges: the need for extensive training data and limited cross-scenario transferability. While some methods achieve strong performance within specific scenarios (Allen et al., , 2020), they often fail to generalize their understanding to novel situations, falling short of human-level reasoning capabilities (Kang et al., , 2025). Our work addresses these limitations by introducing a framework that unifies passive observation and interactive learning through dynamic force fields. The NFF approach enables few-shot learning of physical principles while supporting active reasoning through its interpretable force-based representation, facilitating both accurate prediction and effective intervention planning across diverse physical scenarios.
@@
Dynamic prediction  The prediction of physical dynamics, fundamental to intuitive physics, has evolved along two methodological branches: discrete and continuous-time approaches. Discrete methods typically combine recurrent architectures with GNNs (Battaglia et al., , 2016; Qi et al., , 2021) or transformer modules (Wu et al., , 2023) for modeling object interactions, while leveraging convolutional operators for processing pixel-based information (Shi et al., , 2015; Wang et al., , 2022). Despite their flexibility in handling various interaction types, these methods often struggle with two key limitations: difficulty in extracting robust physical representations and susceptibility to error accumulation over extended time horizons.
@@
Continuous-time methods address the above issues by incorporating physical inductive biases via explicit modeling of state derivatives within dynamical systems (Chen et al., , 2018; Zhong et al., , 2020; Norcliffe et al., , 2020). While they improve long-term prediction, current systems typically assume energy-conservative systems with simplified dynamics (Greydanus et al., , 2019; Cranmer et al., , 2020) or static fields (Kofinas et al., , 2023). More importantly, these methods face significant challenges in few-shot learning and struggle with cross-scenario generalization (Chen et al., , 2020; Yuan et al., , 2024), often requiring specific priors and grammars to achieve meaningful transfer (Xu et al., , 2021). Although Physics-informed Neural Networks (PINN) can be precise in modeling physical dynamics, they depend heavily on prior knowledge of the underlying equations (Chu et al., , 2022).
@@
Our work advances continuous-time methods to model complex rigid body interactions via learnable dynamic force fields. This approach enables robust cross-scenario generalization while handling non-conservative energy systems. By combining the stability advantages of continuous-time modeling with flexible force field representations, NFF achieves both accurate long-term predictions and strong generalization capabilities without requiring explicit physical priors.
@@ 3.1 The Neural Force Field (NFF) representation
Traditional approaches to modeling physical interactions typically rely on implicit representations through hidden vectors to describe object interactions. These methods use neural networks to learn state transition functions that map current scene features to future states. However, this purely data-driven approach cannot guarantee physically grounded representations, leading to poor generalization despite intensive training. We introduce NFF, which addresses these limitations by learning physics-grounded, generalizable representations through explicit force field modeling; see Fig.˜2 for a comparison.
@@
Predicting force field  The core of NFF is built on the physical concept of a force field—a vector field that determines the force experienced by any query object based on its state $\mathbf{z}^{q}$. For a scene with $N$ objects, we model the force field $\mathbf{F}(\cdot)$ at time $t$ using a neural network operating on a relation graph $\mathcal{G}$. The graph contains object nodes $V=\{\mathbf{z}^{0}(t),\mathbf{z}^{1}(t),...,\mathbf{z}^{N-1}(t)\}$. Following neural operator methods (Lu et al., , 2021) and graph neural models (Battaglia et al., , 2018), we formulate the force field function as:
$\small\mathbf{F}(\mathbf{z}^{q}(t))=\sum_{i\in\mathcal{G}(q)}\mathbf{W}\left(f_{\theta}(\mathbf{z}^{i}(t))\odot f_{\phi}(\mathbf{z}^{q}(t))\right)+\mathbf{b},$ (1)
@@
where $\mathcal{G}(q)$ represents the neighbors of query $q$ determined by contacts or attractions, $f_{\theta}$ and $f_{\phi}$ are neural networks with parameters $\theta$ and $\phi$, and $\mathbf{W}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{force}}},\mathbf{b}\in\mathbb{R}^{d_{\text{force}}}$ map hidden features to low-dimensional forces. The state vector $\mathbf{z}$ incorporates zero-order (position, angle) and first-order (velocity, angular velocity) states. This representation can handle various physical interactions including collision, sliding, rotation, and gravity, as shown in Fig.˜1.
Figure 2: Comparison between discrete latent transition in traditional interaction modeling and continuous explicit transition in NFF. Traditional interaction models such as IN and SlotFormer decode discrete frames from a learnable decoder. In contrast, NFF decodes trajectories by integrating the continuous force using an ODE solver, which benefits learning detailed interactions such as collisions.
@@
Decoding trajectory with ODE Rather than relying on neural decoders, NFF employs a second-order ODE integrator to compute object trajectories from the learned force field. This approach ensures physically consistent trajectories governed by fundamental principles. The dynamic change of motion for a query state follows:
$\small\frac{d^{2}x^{q}(t)}{dt^{2}}=\frac{dv^{q}(t)}{dt}\propto\mathbf{F}(\mathbf{z}^{q}(t)).$ (2)
@@
While object mass theoretically affects acceleration, these mass constants are safely absorbed into the learned $F(\cdot)$ either by being explicitly included as inputs or implicitly captured through correlated physical attributes like size.
@@
Solving ODEs We solve the ODE system using integration methods such as Runge-Kutta and Euler:
$\small\mathbf{z}^{q}(t)=\text{ODESolve}\left(\mathbf{z}^{q}(0),\mathbf{F},0,t\right),$ (3)
$\small\begin{dcases}\mathbf{x}(t)=\mathbf{x}(0)+\int_{0}^{t}\mathbf{v}(t)\,dt,\\ \mathbf{v}(t)=\mathbf{v}(0)+\int_{0}^{t}\mathbf{F}(\mathbf{z}^{q}(t))\,dt,\end{dcases}$ (4)
@@
The NFF framework offers three key advantages over existing approaches. First, it enables few-shot learning of force fields from minimal examples—even single trajectories—through its low-dimensional, physically-grounded representation $\mathbf{F}(\mathbf{z}^{q}(t))$, adhering to Eq.˜2. Second, by representing interactions as neural operator force fields, as shown in Eq.˜1, NFF can generalize the learned force patterns to novel scenarios through simple force summation on new interaction graphs, allowing modeling varying object compositions, interaction types, and time horizons. Third, the high-precision integration decoder, formulated in Eqs.˜3 and 4, enables both fine-grained modeling of continuous-time interactions and efficient optimization for planning tasks through its invertible information flow. This precision supports various applications, from determining initial conditions for desired outcomes to designing targeted force fields for specific behaviors. The effectiveness of these three components—ODE grounding, Neural Operator Learning (NOL), and precise integration—is systematically validated through ablation studies in Sec.˜4.5.
@@
Training NFF The framework employs autoregressive prediction, using previously predicted states to generate current states. Network optimization minimizes the Mean Squared Error (MSE) loss between predictions and ground truth. To promote learning of local, generalizable dynamics rather than global patterns, we segment long trajectories into smaller units during training. This alleviates the issue of accumulated error in teacher forcing while reducing the complexity in recursive training. In evaluation, the model predicts all future dynamics given only the initial states.